Tuesday, April 16, 2013, 11am, Construction Trailer
Harold, Dorothy and Bob met from 11:15am to 12 noon.
Stucco color: The new ochre color is being applied today (4/16). We need to wait until it dries before we can tell how it will look. It is good that the final coat is finally going on! The red/rose color should have a result soon too.
Courtyard tiles. The new ones have been delivered and will be installed once the scaffolding has come down. The existing tiles have taken a beating during construction. The project will assess and try to clean up any damage.
Painting non-Project windows blue:
Owner-installed white fiberglass windows need to be sanded, primed and painted by owners (unless Matt van Dyke is doing it, since it is part of his bid), with appropriate products. Bob will try to provide a spec or narrative how the project is doing this.
O’Hagin vents in tower roofs. Note that some towers are marked with an orange “D” (for dormer) to determine whether the vents need to be on the back and front, or whether they can be less visible on the sides. The architect is still working on this. He needs to assure that the same measure of ventilation will be achieved, or calculate the percentage change so we can decide on the trade-off.
Weep Screeds & Permeability. Harold relayed Harry’s concern about the permeability of the Sto coating at the weep screeds. Bob responded that this issue has been addressed by the Sto rep. and Vanderslice, even if no numbers are forthcoming. They would require substantial research time.
Window/door casings: The project will reuse old and replace with new as necessary. They are working on 4 options on how to marry the mullion that comes with the new windows, with the old casing. The architect is involved; this should be resolved today, hopefully with minimal cost impact.
Each solution is being mocked up at Dorothy’s house to make the decision. (We viewed these after the meeting.)
Exterior light fixtures: The order has been placed, arrival date not known yet. The money flow from Board to Project is being handled at the Stonemark office. The bulbs will be replaced with compact fluorescents. (Who will pay?)
Scheduling. As presented at last Saturday’s meeting the completion dates for the 910s is mid-May, for the 900s mid-June and the 920s mid-July. The new deck doors and trellis work for the 920s has been completed (except on door delivered with hinges on the wrong side). Work in the 920s will start again once the door/window crew, now moving unit by unit through the 900s, completes the last unit in that cluster.
Bathroom vent testing/replacement protocol for owners. Stonemark is working on this now, it is almost ready to go out.
Next meeting: Tuesday April 30, 2013, 11am.